OpenStudy (anonymous):

f(p+3)=f^2 can a/o explain how to do this?

OpenStudy (solomonzelman):

factor this equation out fp+3f=f^2 divide this equation by f fp/f+3f/f=f^2/f the “f” s cancel out p+3=f bring 3 to the other side... so the answer is p=f-3

OpenStudy (anonymous):

Tnx